# Real-time Verification and Refinement of Language Model Text Generation

### v4

@@ Abstract
Real-time Verification and Refinement of Language Model Text Generation
Joonho Ko Jinheon Baek Sung Ju Hwang KAIST joon001001@gmail.com, {jinheon.baek, sjhwang82}@kaist.ac.kr
@@ 1 Introduction
Large language models (LLMs) (Achiam et al., 2023; Jiang et al., 2023a; Dubey et al., 2024) have demonstrated significant advancements across various tasks, such as question answering (QA) (Yang et al., 2018; Kwiatkowski et al., 2019; Fan et al., 2019; Min et al., 2020) and its more complex real-world applications supported by information retrieval (IR) (Xiong et al., 2020; Karpukhin et al., 2020; Ni et al., 2021). However, LLMs still face notable limitations like hallucinations, mainly due to the incorrect or outdated knowledges of the model itself (Rawte et al., 2023) and the wrong application and generalization of memorized or retrieved knowledge (Jiang et al., 2024; Xu et al., 2024).
@@
In this work, we propose Streaming-VR (Streaming Verification and Refinement), a method designed to address the issue of error propagation in LLM-generated text. As visually depicted in Figure 1 (b), Streaming-VR evaluates model-generated answers in real-time, identifying the entire sentence and correcting only if its subset is wrong. By employing an external verification model, Streaming-VR verifies errors during the generation process, detects inaccuracies in newly generated sequence of tokens, and promptly corrects them. Because rectification occurs immediately after verification and runs concurrently with text generation, Streaming-VR significantly enhances efficiency and improves the factual accuracy of model outputs. Our experimental results show that when LLMs generate incorrect tokens early in a sequence, it substantially increases the likelihood of subsequent sentences being factually inaccurate. Specifically, approximately 37.6% of the answers in various settings were found to contain factual inaccuracies caused by error propagation (early erroneous tokens), highlighting the critical importance of employing Streaming-VR.
We validate the effectiveness and efficiency of Streaming-VR experimentally on two benchmark datasets: achieving approximately 39.8% and 31.5% higher efficiency for ASQA (Stelmakh et al., 2022) and QuoteSum (Schuster et al., 2024) in average, respectively. We have employed Mistral 7B (Jiang et al., 2023a), LLaMA-3.1 8B (Dubey et al., 2024), and GPT-4o (Achiam et al., 2023).
@@ 2 Related Work
Recent advancements in language models (LMs) (Radford, 2018; Devlin et al., 2019; Liu et al., 2019; Raffel et al., 2023) and LLMs with billions of parameters, have led to significant improvements in performance across various natural language tasks. Since LMs cannot memorize or learn every real-world knowledge, several studies have explored methods to enhance their capabilities by leveraging external knowledge sources like retrieval-augmented generation (Lewis et al., 2020), for knowledge-intensive tasks. Despite the assistance of external knowledge, models often generate incorrect answers due to the failure of factual recall (Jiang et al., 2024) as they may not succeed in retrieving or applying the relevant knowledge appropriately, and generalizing memorized knowledge accurately.
@@
Other than the misinformation induced by wrong knowledge, LLM itself often generates plausible but incorrect texts (Zhang et al., 2024) (i.e., hallucination). Thus, evaluating the factuality (Thorne et al., 2018; Min et al., 2023) of LLM outputs correcting inaccuracies has emerged as an important topic. Various approaches explore methods to enhance the factual accuracy of model responses and develop robust fact-checking or answer-verifying models. For instance, Dhuliawala et al. (2024) generates a series of independent questions to check the factual claims made in the model response, followed by synthesizing the answers from the verification step. Beyond evaluating or verifying the faithfulness of LLM answers, answer-correction has also become a prominent area of focus in various fields. Iterative refinement is well known to be helpful for improving generative contents of natural language (Madaan et al., 2024) and code (Faltings et al., 2021; Yasunaga et al., 2021) autonomously, but limited to the final outcome after waiting for the whole generation is done.
On the other hand, Lightman et al. (2023) demonstrates the effectiveness of process supervision by focusing on each step of the reasoning process, and allowing the model to identify and correct errors in the middle. It emphasizes the importance of intermediate verification in complex multi-step reasoning tasks like mathematical problem solving where a single error can derail the entire answer. Also, Welleck et al. (2023) employs an online training procedure for a separate corrector to learn from feedback on intermediate outputs. Nevertheless, LMs are capable of correcting errors only when their locations are identified (Tyen et al., 2024) exactly, which poses a bottleneck in improving self-correction capabilities. Furthermore, Huang et al. (2024) have demonstrated through experimental analyses that current LLMs struggle to self-correct their reasoning without external feedback, often resulting in degraded performance after attempting self-correction. Alternatively, Cobbe et al. (2021); Wang et al. (2023) utilize a trained critique model or verifier to correct errors on responses through their feedback. In addition, Gou et al. (2024) show that verification and correction can be done effectively by interacting with diverse external tools. In contrast to the previous works which have to wait for the entire answer generation or limited to the inherent answering ability, we propose a novel method with external model, that refines the specific intermediate sentence of an answer identified as incorrect, with higher efficiency.
@@ 3.1 Preliminaries
We begin with preliminaries, formally explaining Large Language Models and traditional verify-and-refine approach, Full Verification and Refinement.
@@
For the real-time sentence-level verification and refinement, we also analyze the individual sentences in answer. To elaborate, an answer is structured as a sequence of sentences, expressed as , where the notation signifies concatenation in the specified order. To facilitate real-time correction of incorrect sentences within intermediate answers, we define the intermediate answer at a certain step () as containing sentences in total. Note that this can also be expressed as , where is the most recently generated sentence in streaming setup. We initialize as an empty string for coherence.
@@
The simplest traditional approach for verifying and refining LLM answers, namely Full-VR (Full Verification and Refinement), is the most common strategy for improving them just by re-generating the entire responses if identified to be incorrect. While many previous works (Yasunaga et al., 2021; Cobbe et al., 2021; Welleck et al., 2023) achieve significant improvements through supplementary techniques, we focus solely on the vanilla setting, for a direct efficiency comparison without any additional methods designed to increase the factual accuracy of answers. And finally, the overall Full-VR pipeline is expressed as follow for a given query and its answer :
@@ 3.2 Streaming Verification and Refinement
Our approach is structured in the following steps during the generation of answers: 1) Streaming-Verification and 2) Streaming-Refinement if necessary; for the sentence identified as error and go back to 1). We formulate the overall framework of Streaming-VR as follow for a given query and the -th sentence in its answer:
where is the verification output, and is the new sentence output of Streaming-VR at a certain step . Note that the refinement model, Refiner takes into account the whole context of previously verified and (may have been) refined sentences, . After processing all the sentences by Streaming-VR, the final refined answer output should be in the form as follow: .
The answer verification relies on the verifier’s output, such that . We utilize a fine-tuned LLM to determine whether the input is True or False by evaluating the factuality of the generated answers in sentence-level. To this end, we augment training data with true- and false-labeled sentences, as there is no proper question answering dataset labeled accurately with unit-level (e.g. sentence-level) answers for our streaming-verifier. The augmented sentences are made from the provided reference answer data by rephrasing it for True and adding wrong information for False by GPT-4o (Achiam et al., 2023) with the specific prompt as in Appendix A. To suit real-time verification scenarios, we split the answer data into individual sentences using NLTK (Bird and Loper, 2004). These sentences are concatenated incrementally in their original order to form intermediate answers , ensuring that False-labeled sentences only appear at the end, never in the middle. This design allows the streaming-verifier to focus on determining the factuality of the newly-generated sentence at the end. To further enhance the training process, a special sentence-separation token, [SEP] is inserted right before the last sentence in each intermediate answer, formatted as for a certain stage . This setup allows a model to be trained to verify the last sentence along with the context from the preceding True-labeled paragraph in the train set.
@@ 4.1 Datasets and Evaluation Metrics
We leverage two different datasets to evaluate the effect of Streaming-VR especially for multi-answer questions which require well-grounded responses to assess the trustworthiness of QA systems.
ASQA (Stelmakh et al., 2022) is a challenging dataset serving as a bridge between factoid and long-form QA tasks by addressing ambiguous questions that can have multiple correct answers depending on their interpretation. It is composed of 4,353 and 948 questions in the train and dev sets, respectively, while the test set is not publicly available. So we use the dev set as our test set here. ASQA provides the reference long-form answers for every questions which are originated from AmbigQA (Min et al., 2020), the ambiguous questions subset of questions from NQ (Kwiatkowski et al., 2019). In this paper, to evaluate the quantitative performance of methods on ASQA, we follow the official metrics and report: Disambiguous-Rouge (DR) as the overall score which combines ROUGE-L (R-L) (Lin, 2004) for text quality and Disambig-F1 (Dis-F1; QA score based on RoBERTa large (Liu et al., 2019)) for factual correctness.
QuoteSum (Schuster et al., 2024) is also a difficult question answering dataset for Semi-Extractive Multi-source Question Answering (SEMQA), a task designed to assess the comprehensive answering ability by summarizing information from multiple sources. To be specific, SEMQA requires models to generate a comprehensive response that integrate verbatim factual spans extracted from input sources along with supplementary non-factual text connecting them, thereby ensuring a cohesive answer. QuoteSum is made up of 4,009 semi-extractive answers to 1,376 unique questions from PAQ (Lewis et al., 2021) and NQ. For the quantitative evaluation on QuoteSum, we follow the official metrics and report: ROUGE-L, Sem-F1 for answer extraction quality, and overall SEMQA score where they do not require any model-based evaluations.
Building on the original evaluation of QuoteSum (Schuster et al., 2024), we further conduct a quantitative assessment of the variants of few-shot models. Specifically, we use dynamic prompt with top- examples for each questions in the test set, as provided in the original paper. These examples are retrieved from the training set by selecting the passages whose queries are most similar to the target test query, based on the cosine similarity between their sentence embeddings (Ni et al., 2022).
@@
4.2 Analyses on Efficiency
In addition to evaluating the quality and factual accuracy of model responses, we also measure token count to assess the efficiency of each method. Since our experiments rely on models accessed through the HuggingFace (Wolf et al., 2020) API, it was not feasible to implement simultaneous execution of the verifier and refiner alongside the answering model, as would occur in real-world applications. Consequently, we analyze the inference cost (i.e., the number of tokens) per model for each method. This metric is crucial as the number of refined tokens directly affects to the LLM user’s waiting time for response corrections. To quantify the efficiency, we define the efficiency of Streaming-VR relative to Full-VR, taking a cue from the thermal efficiency in thermodynamics, which is formulated as: . Here, and represent the average number of generated tokens in the refinement phase per answer for Streaming-VR and Full-VR, respectively.
@@ 4.3 Experimental Results and Analyses
The main results on ASQA and QuoteSum are summarized in Table 1. Both Full-VR and Streaming-VR employ Mistral 7B as the verifier and GPT-4o as the refiner across three different backbone models (Mistral 7B, LLaMA-3.1 8B, and GPT-4o) for answer generation, as indicated in the method column. Across all response models, the final outcomes after verification-and-refinement converge to similar scores, indicating that the overall quality and faithfulness of the answers are largely determined by the refinement model.
However, we observe a notable performance decline when GPT-4o is used as the backbone for answer generation. Both Full-VR and Streaming-VR with GPT-4o lead to significant drops in Disambig-F1 on ASQA, a key metric for assessing the informativeness of long-form answers, and no other improvements on scores of QuoteSum. These results suggest that GPT-4o, which already generated high-quality answers, may be susceptible to over-correction during the refinement process, reducing the overall effectiveness of the responses. This finding highlights a broader trend: refining answers with the same model used for generation—even a powerful model like GPT-4o—may not improve performance and can even degrade it. For the applications like large-scale data analysis or high-frequency user requests handling thousands or millions of queries daily, or individual users requiring detailed lengthy responses, relying on expensive models like GPT-4o for both generation and refinement can quickly exceed budgetary constraints. Therefore, Streaming-VR, which uses a more cost-effective model for response generation and GPT-4o solely for refinement, emerges as a more practical and economical solution.
To assess the consistent efficacy of Streaming-VR across various settings of RAG and ICL for answer generation, we conduct additional experiments as visualized for performance in Figure 2 and for efficiency in Figure 3. The answers are generated by LLaMA-3.1 8B, verified by Mistral 7B and refined by GPT-4o on both datasets. The results show Streaming-VR’s competitive performance compared to Full-VR. Streaming-VR consistently outperforms the initial answers without refinement and achieves comparable results to Full-VR. It also illustrates that Streaming-VR delivers results on par with Full-VR across all retrieved passage and example shot counts, offering performance improvements over the unrefined original response outputs of language model.
@@
Streaming-VR enhances time efficiency too
We further provide a comprehensive analysis of the overall inference costs extending our evaluation beyond token efficiency in Table 2. Streaming-VR consistently produces a significantly smaller number of tokens than Full-VR, as it skips unnecessary modifications to sentences that are already correct. However, beyond the token counts, for practical deployment in real-world applications, latency plays a critical role in assessing efficiency. Specifically, latency is directly influenced by various factors other than token counts, such as the number of model invocations, execution time per call, and whether the models operate in parallel. Compared to the naïve method (purely sequential generation without any verification or refinement), sentence-level correction introduces some inherent delay, as each sentence is verified and refined before proceeding to the next sentence. However, this delay is mitigated by Streaming-VR’s streamlined correction mechanism, where incorrect sentence is processed in parallel with sentence generation and verification. As a result, per-sentence verification and refinement does not accumulate linearly, keeping overall latency manageable. Importantly, the external verifier and refiner in Streaming-VR are invoked only once and they operate in parallel.
We measured the latency of each method by timing the answer refinement process when an answer is determined to be incorrect. Specifically, using Mistral 7B as the verifier and LLaMA-3.1 8B as the refiner, Streaming-VR requires only an average of 3.07s, whereas Full-VR takes an average of 5.98s to refine the entire answer per question. This makes Streaming-VR approximately 1.95 faster than Full-VR while preserving answer factuality. This analysis underscores the novelty of Streaming-VR across the entire pipeline. For more detailed explanation along with mathematical descriptions on general cases, please refer to Appendix C.
@@
The results in Table 3 show that verifier models can be effective without being large. On both tasks, Streaming-MG performs comparably to Streaming-LG, demonstrating that smaller models can still deliver significant performance gains. These findings highlight that the choice of verifiers is very robust in Streaming-VR, leading to choose smaller models that are resource-efficient and effective, making them particularly valuable for real-world applications with limited computational resources.
@@
The results in Table 3 highlight the critical role of a larger and more advanced model for refinement after verification, even when the verifier is relatively small. Using Mistral 7B as verifier and refiner (Streaming-MM) results in no improvement or even degraded performance across datasets and settings.
In contrast, larger refiners yield significant gains. With LLaMA-3.1 8B as the refiner (Streaming-ML), there is a modest Dis-F1 improvement for closed-book setting on ASQA, though handling multiple passages remains challenging. On QuoteSum, Streaming-ML achieves notable improvements in both zero- and five-shot settings, while Streaming-MM reduces answer quality. The most substantial boost comes from GPT-4o as the refiner (Streaming-MG), whose advanced reasoning capabilities drive superior performance in both RAG and ICL settings. These results confirm the importance of using a refiner larger than the response model for coherent and accurate answers.
@@
Additionally, we conduct some experiments to evaluate the efficacy of self-verification and self-refinement within the Streaming-VR pipeline, utilizing only LLaMA-3.1 8B for backbone, verifier and refiner models. In Table 3, the rows of Self-VR (Self-Verification and Refinement; i.e., Streaming-LL) illustrate that LLMs continue to face challenges with intrinsic self-correction with some performance drops. This result strengthens the conclusions drawn by Huang et al. (2024), which also have demonstrated that intrinsic self-correction, an approach that model attempts to rectify its initial responses using only its inherent capabilities without external feedback, degrades the response quality.
@@
As Zhang et al. (2024) point out that the mistakes or hallucinations in the middle of answer can skew the whole response, we report the statistics of model-generated answers with the rate of derailed answers on each dataset. Specifically, the answers are generated by LLaMA-3.1 8B and verified the finetuned streaming verifier as before. The rate of derailed answers is the ratio of the number of ‘answers composed of false sentences in sequence from the first erroneous sentence to the last one’ to ‘false answers if at least one of their sentences is identified as false’. Results in Figure 4 for ASQA and QuoteSum are 26.3% and 48.9% on average across different settings for RAG and ICL, respectively. Therefore, they highlight the importance of Streaming-VR to prevent derailed responses.

### v5

# Efficient Real-time Refinement of Language Model Text Generation

@@ Abstract
Efficient Real-time Refinement of Language Model Text Generation
Joonho Ko Jinheon Baek Sung Ju Hwang KAIST {joonho.ko, jinheon.baek, sungju.hwang}@kaist.ac.kr
@@ 1 Introduction
Large language models (LLMs) (Achiam et al., 2023; Jiang et al., 2023a; Dubey et al., 2024) have demonstrated significant advancements across various tasks, such as question answering (QA) (Yang et al., 2018; Kwiatkowski et al., 2019; Fan et al., 2019; Min et al., 2020) and their more complex real-world applications, which are even supported by information retrieval (IR) for accurately generating answers (Xiong et al., 2020; Karpukhin et al., 2020; Ni et al., 2021). However, LLMs still face notable limitations like hallucinations, mainly due to the incorrect or outdated knowledge of the model itself (Rawte et al., 2023) and the wrong application and generalization of memorized or retrieved knowledge (Jiang et al., 2024; Xu et al., 2024).
@@
In this work, we propose Streaming Verification and Refinement, in short Streaming-VR, which is a method designed to address the issue of error propagation in LLM-generated text. As visually depicted in Figure˜1 (b), Streaming-VR evaluates model-generated answers in real-time, identifying the entire token sequence and correcting only if its subset is wrong. By employing an external verification model, Streaming-VR verifies errors during the generation process, detects inaccuracies in the newly generated sequence of tokens, and promptly corrects them. Because rectification occurs immediately after verification and runs concurrently with text generation, Streaming-VR significantly enhances efficiency and improves the factual accuracy of model outputs. Our experimental results show that when LLMs generate incorrect tokens early in a sequence, it substantially increases the likelihood of subsequent sentences being factually inaccurate. Specifically, approximately 37.6% of the answers in various settings were found to contain factual inaccuracies caused by error propagation (early erroneous tokens), highlighting the critical importance of employing Streaming-VR.
We validate the effectiveness and efficiency of Streaming-VR experimentally on two benchmark datasets: achieving approximately 39.8% and 31.5% higher efficiency for ASQA (Stelmakh et al., 2022) and QuoteSum (Schuster et al., 2024) in average, respectively. Also, Streaming-VR is approximately 1.95 faster than Full-VR with comparable answer quality. We have employed Mistral 7B (Jiang et al., 2023a), LLaMA-3.1 8B (Dubey et al., 2024), and GPT-4o (Achiam et al., 2023).
@@ 2 Related Work
Recent advancements in language models (LMs) (Radford, 2018; Devlin et al., 2019; Liu et al., 2019; Raffel et al., 2023) and LLMs with billions of parameters have led to significant improvements in performance across various natural language tasks. Since LMs cannot memorize or learn every real-world knowledge, several studies have explored methods to enhance their capabilities by leveraging external knowledge sources like retrieval-augmented generation (Lewis et al., 2020), for knowledge-intensive tasks. Despite the assistance of external knowledge, models often generate incorrect answers due to the failure of factual recall (Jiang et al., 2024) as they may not succeed in retrieving or applying the relevant knowledge appropriately, and generalizing memorized knowledge accurately.
@@
Other than the misinformation induced by wrong knowledge, LLM itself often generates plausible but incorrect texts (Zhang et al., 2024) (i.e., hallucination). Thus, evaluating the factuality (Thorne et al., 2018; Min et al., 2023) of LLM outputs correcting inaccuracies has emerged as an important topic. Various approaches explore methods to enhance the factual accuracy of model responses and develop robust fact-checking or answer-verifying models. For instance, Dhuliawala et al. (2024) generates a series of independent questions to check the factual claims made in the model response, followed by synthesizing the answers from the verification step. Beyond evaluating or verifying the faithfulness of LLM answers, answer-correction has also become a prominent area of focus in various fields. Iterative refinement is well known to be helpful for improving generative contents of natural language (Madaan et al., 2024) and code (Faltings et al., 2021; Yasunaga et al., 2021) autonomously, but is limited to the final outcome after waiting for the whole generation to be done.
On the other hand, Lightman et al. (2023) demonstrates the effectiveness of process supervision by focusing on each step of the reasoning process, and allowing the model to identify and correct errors in the middle. It emphasizes the importance of intermediate verification in complex multi-step reasoning tasks like mathematical problem solving, where a single error can derail the entire answer. Also, Welleck et al. (2023) employs an online training procedure for a separate corrector to learn from feedback on intermediate outputs. Nevertheless, LMs are capable of correcting errors only when their locations are identified (Tyen et al., 2024) exactly, which poses a bottleneck in improving self-correction capabilities. Furthermore, Huang et al. (2024) have demonstrated through experimental analyses that current LLMs struggle to self-correct their reasoning without external feedback, often resulting in degraded performance after attempting self-correction. Alternatively, Cobbe et al. (2021); Wang et al. (2023) utilize a trained critique model or verifier to correct errors on responses through their feedback. In addition, Gou et al. (2024) show that verification and correction can be done effectively by interacting with diverse external tools. In contrast to the previous works, which have to wait for the entire answer generation or are limited to the inherent answering ability, we propose a novel method with an external model that refines the specific intermediate sentence of an answer identified as incorrect, with higher efficiency.
@@ 3.1 Preliminaries
We begin with preliminaries, explaining Large Language Models and the traditional verify-and-refine approach, Full Verification and Refinement.
@@
For the real-time sentence-level verification and refinement, we also analyze the individual sentences in the answer. To elaborate, an answer is structured as a sequence of sentences, expressed as , where the notation signifies concatenation in the specified order. To facilitate real-time correction of incorrect sentences within intermediate answers, we define the intermediate answer at a certain step () as containing sentences in total. Note that this can also be expressed as , where is the most recently generated sentence in a streaming setup. We initialize as an empty string for coherence.
@@
The simplest traditional approach for verifying and refining LLM answers, namely Full-VR (Full Verification and Refinement), is the most common strategy for improving them just by re-generating the entire responses if identified as incorrect. While many previous works (Yasunaga et al., 2021; Cobbe et al., 2021; Welleck et al., 2023) achieve significant improvements through supplementary techniques, we focus solely on the vanilla setting, for a direct efficiency comparison without any additional methods designed to increase the factual accuracy of answers. And finally, the overall Full-VR pipeline is expressed as follows for a given query and its answer :
@@ 3.2 Streaming Verification and Refinement
Our approach is structured in the following steps during the generation of answers: 1) Streaming-Verification, and 2) Streaming-Refinement if necessary (for the sentence identified as an error) and then go back to 1). We formulate the overall framework of Streaming-VR for a given query and the -th sentence in its answer, as follows:
where is the verification output, and is the new sentence output of Streaming-VR at a certain step . Note that the refinement model, Refiner takes into account the whole context of previously verified and (may have been) refined sentences, . After processing all the sentences by Streaming-VR, the final refined answer output should be in the form as follows: .
The answer verification relies on the verifier’s output, such that . We utilize a fine-tuned LLM to determine whether the input is True or False by evaluating the factuality of the generated answers at the sentence-level. To this end, we augment training data with true- and false-labeled sentences, as there is no proper question answering dataset labeled accurately with unit-level (e.g. sentence-level) answers for our streaming-verifier. The augmented sentences are made from the provided reference answer data by rephrasing it for True and adding wrong information for False by GPT-4o (Achiam et al., 2023) with the specific prompt as in Implementation Details (Appendix A). To suit real-time verification scenarios, we split the answer data into individual sentences using NLTK (Bird and Loper, 2004). These sentences are concatenated incrementally in their original order to form intermediate answers , ensuring that False-labeled sentences only appear at the end, never in the middle. This design allows the streaming verifier to focus on determining the factuality of the newly-generated sentence at the end. To further enhance the training process, a special sentence-separation token, [SEP], is inserted right before the last sentence in each intermediate answer, formatted as for a certain stage . This setup allows a model to be trained to verify the last sentence along with the context from the preceding True-labeled paragraph in the train set.
@@ 4.1 Datasets and Evaluation Metrics
We use two different datasets to evaluate the effect of Streaming-VR, especially for multi-answer questions, which require well-grounded responses to assess the trustworthiness of QA systems.
@@
ASQA (Stelmakh et al., 2022) is a challenging dataset serving as a bridge between factoid and long-form QA tasks by addressing ambiguous questions that can have multiple correct answers depending on their interpretation. It is composed of 4,353 and 948 questions in the train and dev sets, respectively, while the test set is not publicly available. So we use the dev set as our test set here. ASQA provides the reference long-form answers for every question, which originate from AmbigQA (Min et al., 2020), the ambiguous questions subset of questions from NQ (Kwiatkowski et al., 2019). In this paper, to evaluate the quantitative performance of methods on ASQA, we follow the official metrics and report: Disambiguous-Rouge (DR) as the overall score, which combines ROUGE-L (R-L) (Lin, 2004) for text quality and Disambig-F1 (Dis-F1; QA accuracy score based on RoBERTa large (Liu et al., 2019)) for factual correctness.
QuoteSum (Schuster et al., 2024) is also a difficult question answering dataset for Semi-Extractive Multi-source Question Answering (SEMQA), a task designed to assess the comprehensive answering ability by summarizing information from multiple sources. Specifically, SEMQA requires models to generate a response that integrates verbatim factual spans extracted from input sources along with supplementary non-factual text connecting them, thereby ensuring a cohesive answer. QuoteSum is made up of 4,009 semi-extractive answers to 1,376 unique questions from PAQ (Lewis et al., 2021) and NQ. For the quantitative evaluation on QuoteSum, we follow the official metrics and report: ROUGE-L, Sem-F1 for answer extraction quality, and overall SEMQA score, where they do not require any model-based evaluations.
Building on the original evaluation of QuoteSum (Schuster et al., 2024), we further conduct a quantitative assessment of the variants of few-shot models. Specifically, we use a dynamic prompt with top- examples for each question in the test set, as provided in the original paper. These examples are retrieved from the training set by selecting the passages whose queries are most similar to the target test query, based on the cosine similarity between their sentence embeddings (Ni et al., 2022).
@@
4.2 Analyses on Efficiency
In addition to evaluating the quality and factual accuracy of model responses, we also measure token count to assess the efficiency of each method. Since our experiments rely on models accessed through the HuggingFace (Wolf et al., 2020) API, it was not feasible to implement simultaneous execution of the verifier and refiner alongside the answering model, as would occur in real-world applications. Consequently, we analyze the inference cost (i.e., the number of tokens) per model for each method. This metric is crucial as the number of refined tokens directly affects the LLM user’s waiting time for response corrections. To quantify the efficiency, we define the efficiency of Streaming-VR relative to Full-VR, taking a cue from the thermal efficiency in thermodynamics, which is formulated as: . Here, and represent the average number of generated tokens in the refinement phase per answer for Streaming-VR and Full-VR, respectively.
@@ 4.3 Experimental Results and Analyses
The main results on ASQA and QuoteSum are summarized in Table˜1. Both Full-VR and Streaming-VR employ Mistral 7B as the verifier and GPT-4o as the refiner across three different backbone models (Mistral 7B, LLaMA-3.1 8B, and GPT-4o) for answer generation, as indicated in the method column. Across all response models, the final outcomes after verification-and-refinement converge to similar scores, indicating that the overall quality and faithfulness of the answers are largely determined by the refinement model.
However, we observe a notable performance decline when GPT-4o is used as the backbone for answer generation. Both Full-VR and Streaming-VR with GPT-4o lead to significant drops in Disambig-F1 on ASQA, a key metric for assessing the informativeness of long-form answers, and no other improvements on scores of QuoteSum. These results suggest that GPT-4o, which already generated high-quality answers, may be susceptible to over-correction during the refinement process, reducing the overall effectiveness of the responses. This finding highlights a broader trend: refining answers with the same model used for generation—even a powerful model like GPT-4o—may not improve performance and can even degrade it. For applications like large-scale data analysis or high-frequency user requests handling thousands or millions of queries daily, or individual users requiring detailed, lengthy responses, relying on expensive models like GPT-4o for both generation and refinement can quickly exceed budgetary constraints. Therefore, Streaming-VR, which uses a more cost-effective model for response generation and GPT-4o solely for refinement, emerges as a more practical and economical solution.
To assess the consistent efficacy of Streaming-VR across various settings of RAG and ICL for answer generation, we conduct additional experiments as visualized for performance in Figure˜2 and for efficiency in Figure˜3. The answers are generated by LLaMA-3.1 8B, verified by Mistral 7B and refined by GPT-4o on both datasets. The results show Streaming-VR’s competitive performance compared to Full-VR. Streaming-VR consistently outperforms the initial answers without refinement and achieves comparable results to Full-VR. It also illustrates that Streaming-VR delivers results on par with Full-VR across all retrieved passages and example shot counts, offering performance improvements over the unrefined original response outputs of the language model.
@@
Streaming-VR enhances time efficiency too
We further provide a comprehensive analysis of the overall inference costs extending our evaluation beyond token efficiency in Table˜2. Streaming-VR consistently produces a significantly smaller number of tokens than Full-VR, as it skips unnecessary modifications to sentences that are already correct. However, beyond the token counts, for practical deployment in real-world applications, latency plays a critical role in assessing efficiency. Specifically, latency is directly influenced by various factors other than token counts, such as the number of model invocations, execution time per call, and whether the models operate in parallel. Compared to the naïve method (purely sequential generation without any verification or refinement), sentence-level correction introduces some inherent delay, as each sentence is verified and refined before proceeding to the next sentence. However, this delay is mitigated by Streaming-VR’s streamlined correction mechanism, where the incorrect sentence is processed in parallel with sentence generation and verification. As a result, per-sentence verification and refinement do not accumulate linearly, keeping overall latency manageable. Importantly, the external verifier and refiner in Streaming-VR are invoked only once and operate in parallel.
We measured the latency of each method by timing the answer refinement process when an answer is determined to be incorrect. Specifically, using Mistral 7B as the verifier and LLaMA-3.1 8B as the refiner, Streaming-VR requires only an average of 3.07s, whereas Full-VR takes an average of 5.98s to refine the entire answer per question. This makes Streaming-VR approximately 1.95 faster than Full-VR while preserving answer factuality. This analysis underscores the novelty of Streaming-VR across the entire pipeline. For a more detailed explanation along with mathematical descriptions on general cases, please refer to Appendix C.
@@
The results in Table˜3 show that verifier models can be effective without being large. On both tasks, Streaming-MG performs comparably to Streaming-LG, demonstrating that smaller models can still deliver significant performance gains. These findings highlight that the choice of verifiers is very robust in Streaming-VR, leading to the choice of smaller models that are resource-efficient and effective, making them valuable for real-world applications with limited computational resources.
@@
The results in Table˜3 highlight the critical role of a larger and more advanced model for refinement after verification, even when the verifier is relatively small. Using Mistral 7B as verifier and refiner (Streaming-MM) results in no improvement or even degraded performance across datasets and settings.
In contrast, larger refiners yield significant gains. With LLaMA-3.1 8B as the refiner (Streaming-ML), there is a modest Dis-F1 improvement for the closed-book setting on ASQA, though handling multiple passages remains challenging. On QuoteSum, Streaming-ML achieves notable improvements in both zero- and five-shot settings, while Streaming-MM reduces answer quality. The most substantial boost comes from GPT-4o as the refiner (Streaming-MG), whose advanced reasoning capabilities drive superior performance in both RAG and ICL settings. These results confirm the importance of using a refiner larger than the response model for coherent and accurate answers.
@@
Additionally, we conduct some experiments to evaluate the efficacy of self-verification and self-refinement within the Streaming-VR pipeline, utilizing only LLaMA-3.1 8B for backbone, verifier and refiner models. In Table˜3, the rows of Self-VR (Self-Verification and Refinement; i.e., Streaming-LL) illustrate that LLMs continue to face challenges with intrinsic self-correction with some performance drops. This result strengthens the conclusions drawn by Huang et al. (2024), which also have demonstrated that intrinsic self-correction, an approach that model attempts to rectify its initial responses using only its inherent capabilities without external feedback, degrades the response quality.
@@
As Zhang et al. (2024) point out that the mistakes or hallucinations in the middle of the answer can skew the whole response, we report the statistics of model-generated answers with the rate of derailed answers on each dataset. Specifically, the answers are generated by LLaMA-3.1 8B and verified by the finetuned streaming verifier as before. The rate of derailed answers is the ratio of the number of ‘answers composed of false sentences in sequence from the first erroneous sentence to the last one’ to ‘false answers if at least one of their sentences is identified as false’. Results in Figure˜4 for ASQA and QuoteSum are 26.3% and 48.9% on average across different settings for RAG and ICL, respectively. Therefore, they highlight the importance of Streaming-VR to prevent derailed responses.